OpenStudy (anonymous):

which is an equation in form for the given point and slope. point (-8,3) slope (6)

OpenStudy (jjwalton04):

Well Heres the format, Y=mx+b so just insert the data u have up there

OpenStudy (anonymous):

but which is the point

OpenStudy (jjwalton04):

Whats the choices?

OpenStudy (anonymous):

y+3=6x-48 y+3=6(x+8) y-3=6(x-8) y-3=6(x+8)

OpenStudy (jjwalton04):

Well im thinking its b but, but not 100%

OpenStudy (jjwalton04):

Hold on gonna tag some people to help out

OpenStudy (jjwalton04):

@Mehek14 can u help us out?

OpenStudy (jjwalton04):

@sourwing

OpenStudy (jjwalton04):

@zepdrix

zepdrix (zepdrix):

Hey, heyheyhey :) It looks like your choices are given in point-slope form. This is really convenient since they gave us a point, and a slope value to work with! Here is the information they gave us:\[\large\rm (\color{orangered}{x_1},\color{royalblue}{y_1})=(\color{orangered}{-3},\color{royalblue}{8}), \qquad\qquad \color{green}{m=6}\]And here is the formula for point-slope form of a line:\[\large\rm y-\color{royalblue}{y_1}=\color{green}{m}(x-\color{orangered}{x_1})\]

zepdrix (zepdrix):

Just plug in the stuff. Match up the colors! :)

OpenStudy (anonymous):